Line L has the equation 4y-6x=33. Line M goes through the point A (5,6) and the point B (-4,k). Line L is perpendicular to M. Work out the value of k.


1
Expert's answer
2020-09-21T14:18:13-0400

So the straight line is given by the equation: "y = k * x + b"

Let write line L as: "y_1(x) = k_1 * x + b_1"

Line L: "4 * y - 6 * x = 33"

Let write line L in general form: "y = \dfrac{33 + 6 * x}{4} = 1.5 * x + 8.25"

From L line's equation: "k_1 = 1.5" , "b_1 = 8.25"

Let write line M as: "y_2(x) = k_2 * x + b_2"

Perpendicularity condition of two lines is: "k_2 = - \dfrac{1}{k_1}"

"k_2 = -\dfrac{1}{1.5} = -\dfrac{2}{3}"

Substitute the numbers in M line's equation:

"M: y_2(x) = -\dfrac{2}{3} * x + b_2"

Let put in this equation values from point "A(x_a = 5, y_a = 6)"

"M: 6 = -\dfrac{2}{3} * 5 + b_2"

"b_2=6 + \dfrac{2}{3} * 5=\dfrac{28}{3}"

Substitute "b_2" in M line's equation:

"M: y_2(x) = -\dfrac{2}{3} * x + \dfrac{28}{3}"

Let put in this equation values from point "B(x_b = -4, y_b = k)"

"M: k = -\dfrac{2}{3} * (-4) + \dfrac{28}{3} = \dfrac{8+28}{3}=\dfrac{36}{3}=12"

"k = 12"